OpenStudy (anonymous):

2/x multiplied by what equals 1?

OpenStudy (anonymous):

x/2

OpenStudy (anonymous):

x/2

OpenStudy (anonymous):

2/x * x/2=1

OpenStudy (sasogeek):

x/2

OpenStudy (anonymous):

spam x/2

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

you guys im dumb at math..

OpenStudy (anonymous):

think about it algebraicly: 2/x * y = 1 Solve for y and you get x/2 But as a general rule, multiplying a/b * b/a = 1

OpenStudy (anonymous):

2/x * y = 1 y=x/2

OpenStudy (sasogeek):

you'll get to know math with practice ;)

OpenStudy (sasogeek):

everyone is dumb at first ;) but smart at last... u know why... cos everyone practices :D so should u :)